Juniors boat will go 15miles per hour in still water.If he can go 12miles downstream in the same amount of time as it takes to go 9 miles upstream,what is the speed of the current?
Found 2 solutions by mananth, amalm06:
Answer by mananth(16946)

let speed of current be x mph
speed of boat in still water = 15 mph
..
downstream speed = x+15
upstream speed = 15-x
..
downstream distance = 12 miles
upstream distance = 9 miles.
...
t= d/r
..
upstream time = downstream time
12/(x+15)= 9/(15-x)
12(15-x)=9(x+15)
180-12x=9x+135
21x=180-135
21x=45
/21
x= 45/21
x= 2.14 mph speed of current
..
m.ananth@hotmail.ca

Answer by amalm06(224)

Still water: 15 mph
Downstream: 15+x
Upstream: 15-x
Distance=RateXTime, so Time=Distance/Rate
Equate the two times:
12/(15+x)=9/(15-x)
Solve for x:
180-12x=135+9x
45=21x
x=2.14 mph (Answer)